three charges are at the corners of an equilateral triangle having charge 0.5c,-.25c and 0.1c.find resultant force on 0.1c charge?

Grade:12

2 Answers

Debomoy Roy
6 years ago
Let,  F1 = 0.5 c   ;  F2 = – 0.25 c    ;  F3 = 0.1 
Since the triangle is equilaterial , hence Angle = 60 (degree)
Let, the sides of the triangle is ‘d’.
now, F31= k(0.5 * 0.1)/d = k 0.05/d
& F32 = k(-0.25*0.1)/d = k 0.025/d 
Therefore, | F |2 = (F31)2 + (F32)2 + 2. F31. F32, cos(60)
(F31)2 + (F32)2 +  F31. F32,
= (1.875 * 10-3)k/ d2
=> F  = ( 4.43 * 10-3)k / d
= ( 4.43 * 10-3 * 9 * 109)/d [Since k = 9 * 109 ]
= ( 3.987 * 107) / d. 
This is the resultant force on 0.1 c charge.
Vikas TU
6 years ago
Without distance, no one can find out the resultant force on it.
That is, here q1 = 0.5c , q2 = -0.25 C , q3 = 0.1C
Resultant force will be  = q2*q3/r^2.